The length of a rectangle is 5 inches more than double its width. If the perimeter is 70 inches, find the dimensions of the rectangle by using a system of equations.
Thank you.

Answer by tommyt3rd(5050)

length of a rectangle is 5 inches more than double its width means:
l=2w+5
the perimeter is 70 inches means:
2l+2w=70
or l+w=35
subbing in for l and simplifying...
(2w+5)+w=35
3w=30
so w=10
then that means
l+10=35
and so l=25
the dimensions must be: 25 inches x 10 inches